Find the ratio of the following:
 30 minutes to 1.5 hours

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

30 minutes to 1.5 hours
30 min = 30 / 60
= 0.5 hours
Required ratio = (0.5×1) /(0.5 × 3)
=1 / 3

Thanks